# Can LLMs Really Learn to Translate a Low-Resource Language from One Grammar Book?

### v1

@@ Abstract
Extremely low-resource (XLR) languages lack substantial corpora for training NLP models, motivating the use of all available resources such as dictionaries and grammar books. Machine Translation from One Book (Tanzer et al., 2024) suggests prompting long-context LLMs with one grammar book enables English–Kalamang translation, an unseen XLR language—a noteworthy case of linguistic knowledge helping an NLP task. We investigate whether the book’s grammatical explanations or its parallel examples are most effective for learning XLR translation, finding almost all improvement stems from the parallel examples. Further, we find similar results for Nepali, a seen low-resource language, and achieve performance comparable to an LLM with a grammar book by simply fine-tuning an encoder-decoder translation model. We then investigate where grammar books help by testing two linguistic tasks, grammaticality judgment and gloss prediction, and we explore what kind of grammatical knowledge helps by introducing a typological feature prompt that achieves leading results on these more relevant tasks. We thus emphasise the importance of task-appropriate data for XLR languages: parallel examples for translation, and grammatical data for linguistic tasks. As we find no evidence that long-context LLMs can make effective use of grammatical explanations for XLR translation, we suggest data collection for multilingual XLR tasks such as translation is best focused on parallel data over linguistic description.
@@ 1 Introduction
Most of the world’s languages are extremely low-resource (XLR), severely lacking in suitable corpora for NLP tasks (Joshi et al., 2020; Ranathunga & de Silva, 2022), such as parallel data for machine translation (MT). However, over 50% of languages have both a dictionary and a grammar (Nordhoff & Hammarström, 2011). While human-readable, grammar texts are less easily incorporated into standard NLP models due to their non-standard, unstructured format. Large language models (LLMs), which competently handle free-form textual instructions, provide a potential solution to this data mismatch. After pre-training on trillions of tokens (in mainly high-resource languages), LLMs can learn tasks from only a few in-context examples (Brown et al., 2020; Wei et al., 2022). For this reason, interest in grammar texts for NLP tasks is growing (Tanzer et al., 2024; Zhang et al., 2024a; b).
Machine Translation from One Book (Tanzer et al., 2024) claims LLMs can learn to translate between Kalamang (ISO 639-3: kgv)—a newly-documented language unseen in LLM training data—and English (eng) using only a grammar book in-context. We note that kgv is not unique as an XLR language: it has over 3,000 parallel sentences, a dictionary with over 3,000 definitions (Visser, 2020), a 500-page grammar book (Visser, 2022) consisting of grammatical explanations and over 1000 parallel glossed examples, and nearly 100 typological feature specifications (Skirgård et al., 2023a; b). These resources place kgv alongside thousands of XLR languages (OLAC, 2024), and we expect most are also minimally represented in LLMs’ pretraining data. Given this, finding methods to effectively exploit the available kgv resources could have wider implications for XLR NLP.
In this paper, we question the claimed utility of grammatical explanations for XLR MT with LLMs, then ask where and what kind of grammatical knowledge helps. We show that:
Parallel examples are essential for translation We disentangle grammar books’ parallel examples from grammatical explanations, finding explanations add no significant advantage over parallel data: adding 0.7 ChrF++ into kgv, while into eng the book lags 0.3 points against its parallel sentences; and quality drops up to 8 points with parallel data removed. Our findings generalise to Nepali (npi), a seen low-resource language, where the book’s parallel sentences outperform the full book by 4 ChrF++. LLMs fail to effectively exploit grammatical explanations for translation.
@@
Typological prompting outperforms explanations and helps linguistic tasks We introduce a novel typological feature prompt, and for kgv and npi translation we find our method is more effective than explanations into eng, but not into XLR languages. On kgv grammaticality judgment, our typological prompt improves up to 3% over the book’s 1.2k parallel sentences and 8% over the whole book. For gloss prediction, parallel sentences again beat the book by up to 5.3% morpheme accuracy, and adding typology achieves leading performance on this task. Therefore LLMs can exploit grammar for relevant linguistic tasks—if in a useful form—but not for translation.
Task appropriate-data is therefore essential. In the current paradigm, we suggest that data collection for XLR MT is thus better focused on parallel data over linguistic description, given the advantages in token efficiency, computational cost, and availability.
@@ 2 Related Work
Grammar for extremely low-resource machine translation
Translation of low-resource languages (meaning 100k parallel examples) with LLMs is currently of significant interest (Cahyawijaya et al., 2024; Court & Elsner, 2024; Iyer et al., 2024). Methods include fine-tuning (Xu et al., 2024), dictionary prompting (Ghazvininejad et al., 2023), and retrieval-augmented few-shot prompting (Merx et al., 2024). Alongside advances in long-context LLMs, recent work has introduced grammar information in-context for various tasks: Guo et al. (2024) test a textbook-style prompt with LLM-generated parses, seeing limited gains against parallel sentences; and Zhang et al. (2024a) add a singular syntactic rule to their prompt with small effects. Zhang et al. (2024b) chain morphological analysis, a dictionary and an LLM-summarised grammar book in-context, observing small gains from book passages over a dictionary-only setup. Others meanwhile use grammars with LLMs for data augmentation (Lucas et al., 2024) or as a hybrid rule-based translation system (Coleman et al., 2024).
The recent work Machine Translation from One Book (MTOB) (Tanzer et al., 2024) is central to the question of grammar for LLM-based XLR translation. MTOB introduces a translation test set for the newly documented XLR language Kalamang, plus a grammar book and additional parallel sentences. MTOB suggests long-context LLMs can exploit linguistic knowledge (a grammar book) for an NLP task (XLR translation). This represents a strong contribution to XLR NLP, leveraging underused linguistic resources with state-of-the-art LLMs. However, several issues weaken its claims. The test sets of 50 short, easy examples are potentially too small for making wider generalisations, and the human baseline is somewhat flawed as they may learn from examples at test time; relatedly, Gemini Team et al. (2024) make the uncommon step of using a non-fluent annotator to rate model outputs and their own kgv predictions. Furthermore, using ChrF is less appropriate than ChrF++, which factors in word order and is the de-facto standard in XLR translation (Maillard et al., 2023; Costa-jussà et al., 2024; Edman et al., 2024). The results would benefit from further ablations since the signal from parallel sentences and explanations is not disentangled, and we also miss a strong translation approach to the problem. Finally, we note that the grammar book is not designed for learning the language, but rather describes theoretical linguistic phenomena—which the authors note limits LLMs to a basic competence. MTOB therefore leaves open questions over LLMs’ ability to exploit linguistic information for XLR tasks. In this paper, we tackle these issues by combining the test sets, using automatic ChrF++ scores, disentangling the parallel/non-parallel signal, and testing two better aligned tasks for grammar books.
@@
The incorporation of linguistic information into NLP models is a long-standing question with mixed results (Lakoff, 1978; Raskin, 1985; Uszkoreit, 2009; Opitz et al., 2024). Past work sees gains from incorporating syntactic knowledge into translation models using constituency parses (Currey & Heafield, 2019), grammar supertags (Nădejde et al., 2017), or tree-structured models (Sartran et al., 2022). Typological features are a useful form of linguistic description available for many languages in standardised databases (Dryer & Haspelmath, 2013; Skirgård et al., 2023a; b); features describe language variation in terms of phenomena such as word order, verb tenses, and noun cases. Feature specifications are usually drawn from grammar books by linguists, condensing fine-grained, unprocessed textual descriptions into high-level features. Typological features have been incorporated into NLP models with some success in the form of embeddings (Malaviya et al., 2017; Östling & Tiedemann, 2017) but several studies find minimal positive effects on performance (Ponti et al., 2019; Üstün et al., 2022). To test whether LLMs follow this trend, we construct a novel prompt that uses readily available typological feature specifications for source and target languages, as an in-context and language-invariant method for bridging cross-lingual grammatical differences.
@@
Interlinear Glossed Text (IGT) is a structured, standardised data format for linguistic analysis comprising a morphologically segmented transcription (where morphemes are the smallest units of meaning), an interlinear gloss containing lexical and grammatical information, and a translation (Comrie et al., 2015; Mortensen et al., 2023); a Kalamang example is shown in 2 (Visser, 2022). Glossing is an essential step in language documentation (Ginn et al., 2023) and is designed for trained linguists, rather than language learners. Glosses have been widely applied in NLP tasks, including as a pivot for translation (Zhou et al., 2020), dependency parsing (Georgi et al., 2012), grammar generation (Bender et al., 2014), morphological analysis (Moeller et al., 2020; Shandilya & Palmer, 2023), and linguistic resource development (Beermann et al., 2020; Buchholz et al., 2024). Predicting IGT is therefore a well-motivated task, and segmented IGT is a valuable linguistic resource. IGT prediction is most relevant for XLR languages where it is impactful in assisting annotators for documentation and preservation (Ginn et al., 2023). Prior methods include supervised neural models (Zhao et al., 2020; Girrbach, 2023), or adapting multilingual language models (He et al., 2023; Ginn et al., 2024a; b). Since IGT is costly to generate, past work has scraped it from books (Nordhoff, 2020; Nordhoff & Krämer, 2022); we use this method to extract Kalamang glosses from the grammar book. In this paper, we test gloss prediction to better understand whether LLMs can understand and use grammatical knowledge from the grammar book in-context.
@@ 3.1 Grammar Books for Translation
Our methods are guided by open questions over the use of grammar books for XLR translation. First, we manually filter the grammar books into parallel examples and word pairs, and explanatory, descriptive text, to disentangle the signal from translations and grammatical explanations (see Appendix A for a kgv book extract). This novel ablation is necessary to understand which specific aspects of grammar books are useful for XLR MT. We ask whether LLMs really learn effectively from the grammar explanations, or if most translation supervision stems only from the book’s parallel examples. We combine the directional test sets into a single 100 example test set to improve the generalisability of these results, and evaluate with ChrF++ (Popović, 2017) to take word order into account111We omit human evaluation (Gemini Team et al., 2024, cf.) given the infeasibility of engaging proficient Kalamang speakers. See Appendix H for a small-scale qualitative analysis of several kgv–eng examples.. We also test eng–npi translation, a low-resource language with (we expect) a low data weight in LLMs and an established evaluation set, Flores (Costa-jussà et al., 2024); while not unseen, these experiments broaden our results to seen low-resource languages more generally.222Grammar book splits for both kgv and npi will be made available after review.
@@ 3.3 Typological Feature Prompting
In asking what kind of grammatical knowledge can aid LLMs in XLR tasks, we introduce a method for incorporating typological information into prompts. We extract categorical typological feature specifications from Grambank Skirgård et al. (2023b) for kgv, npi, and eng, and use a rule-based template to construct a prompt containing features for each language and a short explanation. For an example of the prompt format, see Appendix B. Most languages with grammar books have some typological feature specification, since features are distilled by trained annotators from external resources. This method therefore isolates the high-level grammatical tendencies of a language from the specific instantiations of those features (i.e. parallel examples). We hypothesise this method, when combined with the grammar book’s parallel sentences, will at least match the performance of the grammar book. We expect that providing explicit features such as word order rules removes some reasoning requirements for the LLM. Conversely, typological features will not have relevant parallel examples, so some reasoning and retrieval is still required, potentially tempering the advantages.
@@ 3.4 Grammaticality Judgment
To test the LLM’s ability to acquire knowledge and understanding of Kalamang grammar from the book, we introduce a discriminative grammar judgment experiment. We ask the model to choose the original Kalamang test sentence against a modified example, with three successively easier settings: swapping two adjacent words (Swapadj), two random words (Swapran), and shuffling all words (Shuffle). We acknowledge that while we cannot guarantee all corruptions are ungrammatical (since no author speaks Kalamang), we assume the original examples are linguistically unmarked sentences. For all settings we expect a 0-shot model to achieve approximately 50% accuracy, while for high-resource languages we would expect near 100% accuracy. We expect the grammar book to have a greater positive impact in this setting where grammatical knowledge is explicitly rewarded.
@@ 3.5 Interlinear Glossed Text Prediction
To explore another more relevant task for exploiting grammar explanations, we test IGT prediction with the grammar book against few-shot and supervised baselines. This experiment tests whether LLMs can learn grammar from a book to the extent that we see a difference in performance on a grammar-focused task. IGT requires both lexical translation and grammar analysis, without any generation in the language at hand. This makes IGT prediction a more appropriate task to perform from a descriptive, non-didactic grammar text. We argue that IGT prediction accelerates XLR documentation more than translation, and is likely to have more direct impact for both first language (L1) speakers and linguists, not to mention the potential downstream uses e.g. POS tagging and MT. IGT prediction is also a well defined task with strong baselines and clear evaluation metrics (Ginn et al., 2023). Finally, we argue grammar books are intuitively suited to IGT prediction more than translation, because their unique contribution is glossed text, rather than just parallel sentences. We use all available sentences with IGT from Dictionaria as our test set, and for our supervised baselines, we process the grammar book IGT examples into a training and development set. We expect the grammar book to provide marginal gains over raw parallel sentences because the grammar book explicitly explains the glossed examples therein.

### v2

@@ Abstract
Extremely low-resource (XLR) languages lack substantial corpora for training NLP models, motivating the use of all available resources such as dictionaries and grammar books. Machine Translation from One Book (Tanzer et al., 2024) suggests that prompting long-context LLMs with one grammar book enables English–Kalamang translation, an XLR language unseen by LLMs—a noteworthy case of linguistics helping an NLP task. We investigate the source of this translation ability, finding almost all improvements stem from the book’s parallel examples rather than its grammatical explanations. We find similar results for Nepali and Guarani, seen low-resource languages, and we achieve performance comparable to an LLM with a grammar book by simply fine-tuning an encoder-decoder translation model. We then investigate where grammar books help by testing two linguistic tasks, grammaticality judgment and gloss prediction, and we explore what kind of grammatical knowledge helps by introducing a typological feature prompt that achieves leading results on these more relevant tasks. We thus emphasise the importance of task-appropriate data for XLR languages: parallel examples for translation, and grammatical data for linguistic tasks. As we find no evidence that long-context LLMs can make effective use of grammatical explanations for XLR translation, we conclude data collection for multilingual XLR tasks such as translation is best focused on parallel data over linguistic description.
@@ 1 Introduction
Most of the world’s languages are extremely low-resource (XLR), severely lacking in suitable corpora for NLP tasks (Ranathunga & de Silva, 2022), such as parallel data for machine translation (MT). However, over 50% of languages have both a dictionary and a grammar (Nordhoff & Hammarström, 2011). While human-readable, grammar texts are difficult to incorporate into most NLP models due to their non-standard, unstructured format. Large language models (LLMs) can handle free-form textual instructions and provide a potential solution to this data mismatch. After pre-training on trillions of tokens (in mainly high-resource languages), LLMs can learn tasks from only a few in-context examples (Brown et al., 2020; Wei et al., 2022). Given this, interest in exploiting grammar texts in-context for NLP tasks is growing (Ramos et al., 2024; Tanzer et al., 2024; Zhang et al., 2024b).
Machine Translation from One Book (Tanzer et al., 2024) claims LLMs can learn to translate between Kalamang (ISO 639-3: kgv)—a newly-documented language unseen in LLM training data—and English (eng) via in-context learning with only a grammar book. We note that kgv has over 3,000 parallel sentences, a dictionary with over 3,000 definitions (Visser, 2020), a 500-page grammar book (Visser, 2022) consisting of grammatical explanations and over 1000 parallel glossed examples, and nearly 100 typological feature specifications (Skirgård et al., 2023a; b). This level of resources is comparable to or more than thousands of XLR languages have (Joshi et al., 2020; OLAC, 2024), thus we expect most of these are also minimally represented in LLMs’ pretraining data. Given this, finding methods to effectively exploit the available kgv resources could have wide-reaching implications for XLR NLP. In this paper, we question the claimed utility of grammatical explanations for XLR MT with LLMs, then ask where and what kind of grammatical knowledge helps. We show that:
Parallel examples are essential for translation We disentangle grammar books’ parallel examples from grammatical explanations, finding explanations add no significant advantage over parallel data: adding 0.7 ChrF++ into kgv, and into eng scores fall 0.3 points adding explanations to parallel sentences; and quality drops up to 8 points with parallel data removed. Our findings generalise to Nepali (npi) and Guarani (gug), where the book’s parallel sentences outperform the full book by up to 4 ChrF++. LLMs fail to effectively exploit grammatical explanations for translation.
@@
Typological prompting outperforms explanations and helps linguistic tasks We introduce a novel typological feature prompt, and for kgv and npi translation we find our method is more effective than explanations into eng, but not into XLR languages. On kgv grammaticality judgment, our typological prompt improves up to 3% over the book’s 1.2k parallel sentences and 8% over the whole book. For gloss prediction, parallel sentences again beat the book by up to 5.3% morpheme accuracy, and adding typology achieves leading performance on this task. Therefore LLMs can exploit grammar for relevant linguistic tasks—if provided in a useful form—but not for translation.
Task-appropriate data is therefore essential. In the current paradigm, we recommend that data collection for XLR MT is thus better focused on parallel data over linguistic description, given the advantages in token efficiency, computational cost, and availability.
@@ 2 Related Work
Grammar for low-resource machine translation
Translation of low and extremely-low resource languages (here meaning 100k and 10k parallel examples respectively) with LLMs is currently of significant interest (Cahyawijaya et al., 2024; Court & Elsner, 2024; Iyer et al., 2024). Methods include fine-tuning (Xu et al., 2024), dictionary prompting (Ghazvininejad et al., 2023), and retrieval-augmented few-shot prompting (Merx et al., 2024). Alongside advances in long-context LLMs, recent work has introduced grammar information in context for various tasks: Guo et al. (2024) test a textbook-style prompt with LLM-generated parses, seeing limited gains against parallel sentences; and Zhang et al. (2024a) add a singular syntactic rule to their prompt with small effects. Zhang et al. (2024b) chain morphological analysis, a dictionary and an LLM-summarised grammar book in-context, observing small gains from book passages over a dictionary-only setup. Others meanwhile use grammars with LLMs for data augmentation (Lucas et al., 2024) or as a hybrid rule-based translation system (Coleman et al., 2024).
Machine Translation from One Book (MTOB) (Tanzer et al., 2024) introduces a translation test set for the newly documented XLR language Kalamang (thus unseen by LLMs), plus a grammar book and additional parallel sentences. MTOB suggests long-context LLMs can exploit linguistic knowledge (a grammar book) for XLR translation, a potential step forward in leveraging underused resources for XLR languages. However, several issues mean MTOB leaves open questions over LLMs’ ability to exploit linguistic information for XLR tasks. The test sets of 50 short, easy examples are potentially too small for making wider generalisations, and the human baseline is somewhat flawed as they may learn from examples at test time; relatedly, Gemini Team et al. (2024) ask the non-fluent human baseline to rate model outputs and their own kgv predictions, potentially biasing the evaluation. Furthermore, despite ChrF++ being the de-facto standard in XLR translation (Maillard et al., 2023; Costa-jussà et al., 2024; Edman et al., 2024), MTOB uses ChrF which unlike ChrF++ does not factor in word order. MTOB’s results would benefit from further ablations, since the signal from parallel sentences and explanations is not disentangled, nor is a strong translation approach tested. Finally, we note that the kgv grammar book is not designed for language learning, but for describing theoretical linguistic phenomena—which MTOB’s authors note limits LLMs to a basic competence. In this paper, we tackle these issues by combining the test sets, using automatic ChrF++ scores, disentangling the parallel/non-parallel signal, and testing two tasks better aligned for grammar books.
@@
Incorporating linguistic information into NLP models is a long-standing goal with mixed results (Lakoff, 1978; Raskin, 1985; Uszkoreit, 2009; Opitz et al., 2024). Past work sees gains from adding syntactic knowledge into translation models using constituency parses (Currey & Heafield, 2019), grammar supertags (Nădejde et al., 2017), or tree-structured models (Sartran et al., 2022). One useful form of linguistic information is typology, available for many languages in standardised feature databases (Dryer & Haspelmath, 2013; Skirgård et al., 2023a; b); features describe languages in terms of phenomena such as word order rules, verb tenses, and noun cases. Trained linguists condense fine-grained textual descriptions from grammar books into discrete, categorical, and cross-linguistically consistent feature specifications. Typological features have been incorporated into NLP models with some success in the form of embeddings (Malaviya et al., 2017; Östling & Tiedemann, 2017) but several studies find minimal positive effects on performance (Ponti et al., 2019; Üstün et al., 2022). To test whether LLMs follow this trend, we construct a novel prompt that uses readily available typological feature specifications for source and target languages, as an in-context and language-invariant method for bridging cross-lingual grammatical differences.
@@
Language documentation involves describing the underlying grammar of a language given its surface forms (Ginn et al., 2023). A standardised data format for this analysis is Interlinear Glossed Text (IGT), comprising a morphologically segmented transcription (where morphemes are the smallest units of meaning), an aligned interlinear gloss with subword-level lexical and grammatical information, and a sentence-level translation (Comrie et al., 2015; Mortensen et al., 2023); a Kalamang example is shown in Example 2 (Visser, 2022). We note that glossing is designed for trained linguists rather than language learners. Glosses have been widely applied in NLP tasks, including as a pivot for translation (Zhou et al., 2020), dependency parsing (Georgi et al., 2012), grammar generation (Bender et al., 2014), morphological analysis (Moeller et al., 2020; Shandilya & Palmer, 2023), and linguistic resource development (Beermann et al., 2020; Buchholz et al., 2024). Predicting IGT is therefore a well-motivated grammatical task, and segmented IGT is a valuable linguistic resource. IGT prediction is most relevant for XLR languages where it is impactful in assisting annotators for documentation and preservation (Ginn et al., 2023). Prior methods include supervised neural models (Zhao et al., 2020; Girrbach, 2023), or adapting multilingual language models (He et al., 2023; Ginn et al., 2024a; b). Since IGT is costly to generate, past work has scraped it from books (Nordhoff, 2020; Nordhoff & Krämer, 2022); we follow this method to extract Kalamang glosses from the grammar book. One of our contributions involves testing gloss prediction to determine whether LLMs can use grammatical knowledge for more relevant tasks.
@@ 3.1 Grammar Books for Translation
Our methods are guided by open questions over the use of grammar books for XLR translation. First, we manually filter the grammar books into parallel examples and word pairs, and explanatory, descriptive text, to disentangle the signal from translations and grammatical explanations (see Appendix A for a kgv book extract). This novel ablation is necessary to understand which specific aspects of grammar books are useful for XLR MT. We ask whether LLMs really learn effectively from the grammar explanations, or if most translation supervision stems only from the book’s parallel examples. We combine the directional test sets into a single 100 example test set to improve the generalisability of these results, and evaluate with ChrF++ (Popović, 2017) to take word order into account111We omit human evaluation (Gemini Team et al., 2024, cf.) given the infeasibility of engaging proficient Kalamang speakers. See Appendix I for a small-scale qualitative analysis of several kgv–eng examples.. We also test eng–npi and gug translation, low-resource languages with an established evaluation set, Flores (Costa-jussà et al., 2024) and likely a low data weight in LLMs; while not unseen, these experiments broaden our results to seen low-resource languages more generally.
@@ 3.3 Typological Feature Prompting
In asking what kind of grammatical knowledge can aid LLMs in XLR tasks, we introduce a text-based method for incorporating typological information into prompts, differing from previous work on continuous typological embeddings (Oncevay et al., 2020). We extract categorical typological feature specifications from Grambank Skirgård et al. (2023b) for kgv, npi, gug, and eng, and use a rule-based template to construct a prompt containing features for each language and a short explanation. For an example of the prompt format, see Appendix D. Most languages with grammar books have some typological feature specification, since features are distilled by annotators from external resources. Our method isolates high-level grammatical tendencies of a language from the specific instantiations of those features (i.e. parallel examples). We hypothesise that our method, when combined with the grammar book’s parallel sentences, will at least match the performance of the grammar book. We expect that providing explicit features such as word order rules removes some reasoning requirements for the LLM. Conversely, typological features will not have relevant parallel examples, so some reasoning and retrieval is still required, potentially tempering the advantages.
@@ 3.4 Grammaticality Judgment
To test the LLM’s ability to acquire knowledge and understanding of Kalamang grammar from the book, we introduce a discriminative grammar judgment experiment. We ask the model to choose the original Kalamang test sentence against a modified example, with three successively easier settings: swapping two adjacent words (Swap), two random words (Swap), and shuffling all words (Shuffle). We acknowledge that while we cannot guarantee all corruptions are ungrammatical (since no author speaks Kalamang), we assume the uncorrupted examples are linguistically unmarked sentences. For all settings we expect a 0-shot model to achieve approximately 50% accuracy, while for high-resource languages we would expect near 100% accuracy. We expect the grammar book to have a greater positive impact in this setting where grammatical knowledge is explicitly rewarded.
@@ 3.5 Interlinear Glossed Text Prediction
To explore another more relevant task for exploiting grammar explanations, we test IGT prediction with the grammar book against few-shot and supervised baselines. This experiment tests whether LLMs can learn grammar from a book to the extent that we see a difference in performance on a grammar-focused task. IGT requires both lexical translation and grammar analysis, without any generation in the language at hand. This makes IGT prediction a more appropriate task to perform from a descriptive, non-didactic grammar text. We argue that IGT prediction accelerates XLR documentation more than translation, and is likely to have more direct impact for both first language (L1) speakers and linguists, not to mention the potential downstream uses, e.g. POS tagging and MT. IGT prediction is also a well defined task with strong baselines from a shared task (Ginn et al., 2023) and clear evaluation metrics, primarily morpheme accuracy (McMillan-Major, 2020; Zhao et al., 2020); our experiments build on this prior work. Finally, we argue grammar books are intuitively suited to IGT prediction more than translation, because their unique contribution is glossed text, rather than just parallel sentences. We use all available sentences with IGT from Dictionaria as our test set, and for our supervised baselines, we process the grammar book IGT examples into a training and development set. We expect the grammar book to provide marginal gains over raw parallel sentences because the grammar book explicitly explains the glossed examples therein.